OpenStudy (anonymous):

last one: domain and range of : 2x^2+y^2=1

OpenStudy (anonymous):

thought u did one like this earlier

OpenStudy (anonymous):

we talked about domain and range, but i got it wrong...

OpenStudy (anonymous):

okay, lemme try

OpenStudy (anonymous):

domain: solve for y right?

OpenStudy (anonymous):

Right..

OpenStudy (anonymous):

okay then: \[y ^{2}= -2x ^{2}+1\] square root both side?

OpenStudy (anonymous):

Yep go ahead..

OpenStudy (anonymous):

okay \[y=\sqrt{-2x ^{2} +1}\] but thats not right since we cant have negatives under the root

OpenStudy (anonymous):

It is right..

OpenStudy (anonymous):

But to find domain it is not right..

OpenStudy (anonymous):

what will you do to find domain here?

OpenStudy (anonymous):

@bronzegoddess I told you the case of square roots.. What we do in case of square roots ??

OpenStudy (anonymous):

greater than or equal to zero

OpenStudy (anonymous):

Yes then go for it..

OpenStudy (anonymous):

so our domain is [1,infinity[ because -2(0)+1= root of 1 which is 1

OpenStudy (anonymous):

\[1 - 2x^2 \ge 0 \implies 2x^2 \le 1 \implies x^2 \le \frac{1}{2} \implies x \le \frac{1}{\sqrt{2}}\]

OpenStudy (anonymous):

ah i should equate to greater or equal to zero....

OpenStudy (anonymous):

If domain will be [1, infinity) then put x = 2 You will get negative that is not right..

OpenStudy (anonymous):

Both..

OpenStudy (anonymous):

you will equate that to greater as well as equal to..

OpenStudy (anonymous):

yes, thats what i meant, so the range will be from our result to infinity?

OpenStudy (anonymous):

Find x in term of y..

OpenStudy (anonymous):

\[Domain : (-\infty, \frac{1}{\sqrt{2}}]\]

OpenStudy (anonymous):

i meant thats our domain?

OpenStudy (anonymous):

oh yes because 2 gives a negative

OpenStudy (anonymous):

Yes...

OpenStudy (anonymous):

Yes..

OpenStudy (anonymous):

so i do the same for x=

OpenStudy (anonymous):

Yes go ahead...

OpenStudy (anonymous):

ok: \[2x ^{2}= 1-y ^{2}\] divide by 2: \[x ^{2}= \frac{ 1-y ^{2} }{ 2 }\] square root: \[x=\sqrt{\frac{ 1-y ^{2} }{ 2 }}\] \[\frac{ 1-y ^{2} }{ 2 } \ge 0\] solve for y:

OpenStudy (anonymous):

@waterineyes shouldn't it be -1/sqrt(2)<=x<=1/sqrt(2)?

OpenStudy (anonymous):

good so far?

OpenStudy (anonymous):

@ghass1978 where?

OpenStudy (anonymous):

Yep My mistake..

OpenStudy (anonymous):

good:)

OpenStudy (anonymous):

Because if you put x = -3 there you will get again negative there..

OpenStudy (anonymous):

am confused

OpenStudy (anonymous):

\[x^2 \le \frac{1}{2} \implies x \le \pm \frac{1}{\sqrt{2}}\]

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

This is assigning a domain a range of : \[[-\frac{1}{\sqrt{2} }, \frac{1}{\sqrt{2}}]\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Yes go ahead @bronzegoddess you are right till there...

OpenStudy (anonymous):

ok ty (: \[-y ^{2}\ge-1 ----> y ^{2}\ge1 --->y \ge \pm1\]

OpenStudy (anonymous):

There is one mistake..

OpenStudy (anonymous):

See when you multiply -1 with an inequality then sign gets reversed..

OpenStudy (anonymous):

where? the

OpenStudy (anonymous):

ah, ok

OpenStudy (anonymous):

\[-y^2 \ge -1\]

OpenStudy (anonymous):

\[y \le \pm1\]

OpenStudy (anonymous):

Here when you multiply -1 then sign reverses.. Yep..

OpenStudy (anonymous):

so range is -1 to 1

OpenStudy (anonymous):

or 1 to -infinity

OpenStudy (anonymous):

See if you have any doubt or confusion then you can verify it.. See you said 1 to - infinity.. Choose one number in this range and plug in for y: For example I choose -2 : \[x = \sqrt{\frac{1 - y^2}{2}}\]

OpenStudy (anonymous):

oh yeah, thats wrong

OpenStudy (anonymous):

\[x = \sqrt{\frac{1 - 4}{2}} = \sqrt{negative}\]

OpenStudy (anonymous):

thx for all the patience and help :)

OpenStudy (anonymous):

So range will be: \[Range : [-1, 1]\]

OpenStudy (anonymous):

It is okay.. Welcome dear.. And I am a boy..